OpenStudy (anonymous):

what is the solution set for x^2-8x+4>-3? x<1 or x>7 71 x>1 or x<7 7

hero (hero):

x^2 - 8x + 4 > - 3 Add 3 to both sides: x^2 - 8x + 4 + 3 > 0 x^2 - 8x + 7 > 0

hero (hero):

The left side factors to (x - 7)(x - 1) > 0 Find the value of the vertex \[\left(-\frac{b}{2a}, f(-\frac{b}{2a})\right)\] Then graph x^2 - 8x + 7

hero (hero):

From there, you can easily observe the parts where the graph is greater than zero

OpenStudy (anonymous):

Thank you ( : does this work for all problems like this?

OpenStudy (anonymous):

what is the variable f? @Hero

hero (hero):

For this set of problems, yes.

hero (hero):

We can refer to the quadratic as f(x) = x^2 - 8x - 7

hero (hero):

Once you find the x coordinate of the vertex, you substitute it in for x, then evaluate

hero (hero):

Have you found the x coordinate for the vertex yet?